OpenStudy (anonymous):

A rectangle has 225 square cm of area. What dimensions result in the least perimeter? (calculus)

OpenStudy (shane_b):

For the calculus method you can do this: http://www.algebra.com/algebra/homework/word/geometry/Geometry_Word_Problems.faq.question.317038.html In reality, I'd just do this:\[L_{side}=\sqrt{225}=15cm\]\[P=4L=4(15cm)=60cm\]

OpenStudy (fellowroot):

Find least perimeter. What do we want? P=2L +2W What do we have? A=LW A=225 225=LW Plug what we have into what we want. W=225/L P=2L+2(225)/L P=2L+450/L To max or min, take derivative, then set derivative equal to zero. dP/dL =2 -450/L^2 0=2-450/L^2 0=2L^2-450 For L Squrt(225) = L 15 =L Now remember.... A=LW=225, L=15 15W=225 W=15 L and W are a min when L=W=15 ....this problem was solved by Eric S. badass math and physics major who will get his phd!